# Bhopal 462 066, Madhya Pradesh, India

### v1

@@ Appendix A Elastic Softening
Appendix B Local Stretching and Tension
@@ Appendix C Conformational Fluctuations
Appendix D Elastic Modulus from the analytical theory
@@
Appendix E IDPs
@@
To understand the influence of charge patterning, we further investigate naturally occurring IDPs namely LAF1 and DDX4, consisting of $N=168$, and $N=236$ residues, respectively79. For these naturally occurring protein variants, the fraction of charged residues (FCR) and the net charge per residue (NCPR) are $0.262$ and $0.024$ for LAF1, and $0.288$ and 0.017-0.017 for DDX4, respectively. The charge pattern of each sequence is quantified using the normalized sequence charge decoration (nSCD), which compares each protein’s charge distribution to the maximum and minimum SCD values possible for its composition93, 61.
$SCD=\frac{1}{N}\sum_{i=2}^{N}\sum_{j=1}^{i-1}q_{i}q_{j}\sqrt{i-j},$ (31)
where $q_{i}$ and $q_{j}$ are charges of the $i^{th}$ and $j^{th}$ monomers, respectively. The nSCD parameter ranges from 0 to 1, where lower values correspond to well-mixed (alternating) charge sequences and higher values indicate charge-segregated (diblock) sequence93, 61.
To model these IDP variants, we employ a modified Lennard–Jones (LJ) potential for non-bonded monomer interactions. This formulation allows the attractive interactions between $i$ and $j$ residues to vary independently from the short-range repulsion, scaled by their average hydropathy $\lambda=(\lambda_{i}+\lambda_{j})/2$94, 95.
@@
The hydropathy $\lambda$ values are based on the Urry scale 96, and the pair potential $U(r)$ is truncated and shifted to zero at a distance of $4\sigma$.

### v2

@@ Acknowledgments
SI-B Local Stretching and Tension
@@
SI-D Elastic Modulus from the analytical theory
@@
SI-E IDPs
To understand the influence of charge patterning, we further investigate naturally occurring IDPs namely LAF1 and DDX4, consisting of $N=168$, and $N=236$ residues, respectively[79]. For these naturally occurring protein variants, the fraction of charged residues (FCR) and the net charge per residue (NCPR) are $0.262$ and $0.024$ for LAF1, and $0.288$ and 0.017-0.017 for DDX4, respectively. The charge pattern of each sequence is quantified using the normalized sequence charge decoration (nSCD), which compares each protein’s charge distribution to the maximum and minimum SCD values possible for its composition[93, 61].
$SCD=\frac{1}{N}\sum_{i=2}^{N}\sum_{j=1}^{i-1}q_{i}q_{j}\sqrt{i-j},$ (31)
where $q_{i}$ and $q_{j}$ are charges of the $i^{th}$ and $j^{th}$ monomers, respectively. The nSCD parameter ranges from 0 to 1, where lower values correspond to well-mixed (alternating) charge sequences and higher values indicate charge-segregated (diblock) sequence[93, 61].
@@
To model these IDP variants, we employ a modified Lennard–Jones (LJ) potential for non-bonded monomer interactions. This formulation allows the attractive interactions between $i$ and $j$ residues to vary independently from the short-range repulsion, scaled by their average hydropathy $\lambda=(\lambda_{i}+\lambda_{j})/2$[94, 95].
@@
The hydropathy $\lambda$ values are based on the Urry scale [96], and the pair potential $U(r)$ is truncated and shifted to zero at a distance of $4\sigma$.